OpenStudy (hulkart):

For the function given, state the starting point for a sample period: f(t)=3cos(pit−1)

OpenStudy (anonymous):

put \[\pi t-1=0\] solve for \(t\)

OpenStudy (anonymous):

then put \[\pi t-1=2\pi\] and solve for \(t\)

OpenStudy (anonymous):

as far as i know there is no such thing as a "starting point" for a period if a function is periodic with period \(p\) it means \[f(x)=f(x+p)\] for all \(x\) there is no "starting point" or "ending point"

OpenStudy (hulkart):

could you give me the answer, pls?